# Overcoming data challenges to measure whole-person health in electronic health records

### v2

@@ Abstract
Objective: The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning. Data from electronic health records (EHR) present a huge opportunity to operationalize the ALI in the learning health system, except they are prone to missingness and errors. Validation of EHR data (e.g., through chart reviews) can provide better-quality data, but realistically, only a subset of patients’ data can be validated, and most protocols do not recover missing data.
Methods: Using a representative sample of $1000$ patients from the EHR at an extensive learning health system ($100$ of whom could be validated), we propose methods to design, conduct, and analyze statistically efficient and robust studies of the ALI and healthcare utilization. With semiparametric maximum likelihood estimation methods, we robustly incorporate all available patient information (validated or unvalidated) into statistical models. Using targeted design strategies, we examine ways to select the most informative patients for validation. Incorporating clinical expertise, we devise a novel validation protocol to promote the quality and completeness of EHR data.
Results: Validation of the EHR data uncovered relatively low error rates and recovered some missing data through auxiliary information in patients’ charts. Through bespoke simulation studies based on preliminary data, residual sampling was identified as the most informative strategy for completing our validation study. Statistical models incorporating the partial validation data indicated higher odds of engaging in the healthcare system were associated with worse whole-person health (i.e., higher ALI), adjusting for age.
Conclusion: Targeted validation with an enriched protocol allowed us to ensure the quality and promote the completeness of the EHR. Findings from our validation study were incorporated into analyses as we operationalize the ALI as a whole-person health measure intended to predict healthcare utilization in the academic learning health system.
Keywords chart review  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ outcome-dependent sampling  $\cdot$ two-phase designs
@@ 1.1 Measuring Whole-Person Health
Learning healthcare systems are transitioning from a reactive (treatment) to a proactive (prevention) philosophy. This transition could prevent future healthcare utilization through early intervention, improving health for the individual and providing cost savings to the institution. Having a standardized summary measure of health that is easy to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
Numerous data types exist within the electronic health record (EHR), including vital signs, demographics, diagnoses, procedures, and clinical notes. Easy access to a standardized and valid summary measure of whole-person health in the EHR would be advantageous for patients, clinicians, and researchers [1, 2]. The electronic frailty index serves this purpose but is only designed for those aged 65 and older [3]. The availability of a whole-person health metric for the general adult population (age $\geq 18$) would allow for longitudinal tracking of individual and population health within a healthcare system, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
Though individual physiological variables like systolic blood pressure measure aspects of whole-person health, they do not provide a comprehensive assessment across various organ systems. The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [4, 5, 6]. The ALI has been used to predict numerous health outcomes (e.g., [7, 8, 9, 10]) and measure racial disparities in health [11]. Because it is sensitive to short-term interventions, the ALI is also an appealing “surrogate” outcome to test health interventions, like lifestyle and environmental modifications.
Many ALI formulations exist. We focus on the calculation from Seeman et al. (2001) [5], which computes the ALI from ten component measurements across three body systems. As illustrated in Supplemental Figure S1, the original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically driven thresholds to define binary “stressor” indicators (e.g., $=1$ if hemoglobin A1c $\geq 6.5$ and $=0$ if hemoglobin A1c $<6.5$). The ALI is ultimately calculated as the sum of these ten stressors, i.e., the number of “unhealthy” measurements the patient had, with values of zero and ten representing the best and worst whole-person health, respectively.
@@ 1.2 Data Quality Challenges in Electronic Health Records Data
EHR data present a huge opportunity to operationalize the ALI in the learning health system. The volume of routinely collected data in the EHR is steadily climbing, increasing the accessibility to clinically meaningful variables. Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains, including HIV and AIDS [12], genetics [13], emergency medicine [14], and therapeutic effectiveness [15].
@@
However, despite clear advantages, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors. These quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [16, 17, 18, 19]. In calculating the ALI, the original measurements may be (i) missing, making the stressor missing, or (i) non-missing but mismeasured, potentially misclassifying the stressor.
@@
Consider a patient whose true underlying HbA1c was $6.5\%$, such that their corresponding stressor should equal one. We illustrate two scenarios for this patient’s data in Figure 1. In the first, they had HbA1c measured, but the value was erroneously transposed to $5.6\%$. Based on the EHR, this patient’s stressor for HbA1c would be misclassified as zero since $5.6\%$ falls into the healthy range. In the second, they did not have HbA1c measured, so the value is missing, and we cannot define a stressor. Simply treating missing components as zero would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining the ALI as the proportion of non-missing stressors a patient experienced.
@@
Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation. In the top scenario, the patient had HbA1c measured but with error, and the correct value was recovered through validation. In the bottom scenario, the patient did not have HbA1c measured, but auxiliary information was located through validation to replace the missing component.
@@
Ignoring errors and missingness in the EHR can hinder analyses, as biased estimates may lead to invalid inference and incorrect conclusions [20, 21]. Errors in EHR data can also reduce the accuracy of prediction models [22]. There is a particular concern for missingness and errors in derived variables, like the ALI, that combine data from many different variables [23, 24]. Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about the association between ALI and healthcare utilization.
@@
For the ALI to be useful in research, it must be reliably and accurately measured. The original calculations, like Seeman et al. (2001) [5], were done in the context of clinical trials, where all patients must have all measurements and data validation protocols are likely in place to ensure quality. Similarly, validation, wherein the extracted EHR data are compared to original source documents (like the electronic chart) provides insights into data quality. Complete data validation (i.e., for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
@@
Fortunately, partially validating EHR data can help identify issues and pave the way for corrections. It provides a cost-effective alternative to complete data validation and is a promising approach to gauge the quality of observational data sources [25, 26, 27]. These partial validation studies are a special case of two-phase designs [28, 29] (Figure 2A). Phase I includes variables available for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation in Phase II.
Figure 2: Diagram of two-phase designs with $N$ patients extracted from the electronic health records (EHR) in phase I and $n$ patients chosen for validation via chart review in phase II. The least informative strategy to select the validation study is A) simple random sampling. More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model. Within strata, the darker shaded areas indicate which patients would be selected.
@@
Validated data are usually expected to be more correct (i.e., less error-prone) but not more complete (i.e., less missingness). To calculate the ALI from EHR data, we needed validation to do both. Herein, we devise a novel validation protocol that incorporates a “roadmap” created by clinical experts to recover missing data created by clinical experts using auxiliary health information (e.g., related diagnoses).
@@
Table 1: Ten component stressors of the allostatic load index were defined by discretizing measurements across three body systems at clinically driven thresholds. If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic. If the auxiliary information was present, the component was treated as a “yes” and included in the patient’s validated ALI.
@@ 1.3 Overview
While very insightful, partial validation studies are expensive, time-intensive undertakings. Using a representative sample of $1000$ patients from the EHR at an extensive academic learning health system ($100$ of whom could be validated), we propose novel methods to design, conduct, and analyze statistically efficient and robust studies of whole-person health (ALI) and healthcare utilization. Using advanced statistics and bioinformatics methods, we seize three key opportunities to make the most of our investment: (i) robust and efficient statistical modeling, (ii) enriched validation protocol, and (iii) targeted validation study design.
@@
Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in the EHR to make patient- and institution-level decisions. However, routinely collected data are prone to missingness and errors.
@@ 2 Methods
2.1 Model and Data
This study was approved by the Institutional Review Board at Wake Forest University School of Medicine. We are interested in modeling healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period. The exposure of interest is the ALI ($X\in[0,1]$), defined as the proportion of stressors a patient experienced. Further, we control for patient age at first encounter ($Z\in\mathcal{R}$).
These variables are assumed to be related through the usual logistic regression model for the log-odds that $Y=1$ (engaged in the healthcare system) given $X$ and $Z$. A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (ORs), denoted ${\boldsymbol{\beta}}$. Due to missingness and errors in the EHR, only an error-prone surrogate measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available. We compute $X^{*}$ as the proportion of non-missing stressors the patient experienced, according to the EHR data. Age $Z$ is fully observed and expected to be error-free.
To obtain consistent estimates for the desired healthcare utilization model, a two-phase design is planned (Figure 2A). In Phase I, the outcome $Y$ and surrogate exposure $X^{*}$ will be collected from the EHR for all $N$ patients. In Phase II, these data will be augmented with the validated exposure $X$ from a chart review for a subset of $n$ patients ($n<N$). To distinguish validated and unvalidated patients, let $V_{i}$ be a validation indicator, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise.
Incorporating data from Phases I and II, along with the validation indicator, the joint distribution of a complete observation is assumed to be
@@ 2.1 Model and Data
where $\Pr(V\mid Y,X^{*},Z)$ is the validation sampling probability; ${\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)$ corresponds to the logistic regression model; $\Pr(X\mid X^{*},Z)$ is the exposure error mechanism; and $\Pr(X^{*},Z)$ is the joint probability density or mass function of $X^{*}$ and $Z$. Two assumptions are made in defining (1). First, selection for validation (i.e., $V$) is assumed to depend only on Phase I variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [30]. Second, the exposure $X^{*}$ from the EHR is assumed to be a surrogate for the validated exposure $X$, such that $Y$ and $X^{*}$ are conditionally independent given $X$. In other words, once $X$ is known, $X^{*}$ is assumed not to contribute anything further about the relationship with $Y$. While other ways of factoring (1) are possible, this factorization avoids making an additional assumption about $\Pr(X^{*},Z)$.
2.2 Robust Statistical Model for Partially Validated Data
The observed-data log-likelihood function (hereafter, log-likelihood function) for the log ORs ${\boldsymbol{\beta}}$ is defined based on the joint distribution in (1), but first, two simplifications can be made. First, the validation sampling probability $\Pr(V|Y,X^{*},Z)$ can be omitted from the log-likelihood for ${\boldsymbol{\beta}}$ because the Phase II variables are MAR. Second, the joint probability $\Pr(X^{*},Z)$ can be ignored because it will drop out as the log-likelihood function is maximized with respect to ${\boldsymbol{\beta}}$ since $X^{*}$ and $Z$ are fully observed.
Observations for all $N$ patients in the study are assumed to be independent and identically distributed (i.i.d.) following (1). Now, the log-likelihood function for ${\boldsymbol{\beta}}$ is proportional to the following:
@@ 2.2 Robust Statistical Model for Partially Validated Data
Notice that unvalidated patients contribute their incomplete information to this log-likelihood by summing over the missing validated exposure $X$:
@@
Still, a model for the exposure error mechanism $\Pr(X\mid X^{*},Z)$ is needed to estimate the analysis model parameters since $X$ is only observed for the subset of validated patients. Since the exposure $X$, representing the patient’s ALI, is bounded between zero and one (inclusive), there are not many flexible, parametric choices to model the distribution of $X$ given ($X^{*},Z$). Thus, a nonparametric estimator is preferable, allowing the validated exposure $X$ to take on any values and arbitrarily depend on $X^{*}$ and $Z$.
@@
Lotspeich et al. (2022) [31] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described. The current setting is a special case with only exposure error, and a brief outline is provided in the Supplemental Materials. For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see Lotspeich et al. (2022) [31].
2.3 Targeted Designs to Select Informative Patients for Validation

### v4

# Overcoming data challenges through enriched validation and targeted sampling to measure whole-person health in electronic health records

@@ Abstract
Objective: The allostatic load index (ALI) is a 10-component composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning. Data from electronic health records (EHR) present a huge opportunity to operationalize the ALI in learning health systems; however, these data are prone to missingness and errors. Validation (e.g., through chart reviews) can provide better-quality data, but realistically, only a subset of patients’ data can be validated, and most protocols do not recover missing data.
Methods: Using a representative sample of $1000$ patients from the EHR at an extensive learning health system ($100$ of whom could be validated), we propose methods to design, conduct, and analyze statistically efficient and robust studies of ALI and healthcare utilization. Employing semiparametric maximum likelihood estimation, we robustly incorporate all available patient information into statistical models. Using targeted design strategies, we examine ways to select the most informative patients for validation. Incorporating clinical expertise, we devise a novel validation protocol to promote EHR data quality and completeness.
Results: Chart reviews uncovered few errors ($99\%$ matched source documents) and recovered some missing data through auxiliary information in patients’ charts. On average, validation increased the number of non-missing ALI components per patient from $6$ to $7$. Through simulations based on preliminary data, residual sampling was identified as the most informative strategy for completing our validation study. Incorporating validation data, statistical models indicated that worse whole-person health (higher ALI) was associated with higher odds of engaging in the healthcare system, adjusting for age.
Conclusion: Targeted validation with an enriched protocol enabled us to ensure the quality and promote the completeness of the EHR. Findings from our validation study were incorporated into analyses as we operationalize the ALI as a scalable whole-person health measure predictive of healthcare utilization in the learning health system.
Keywords chart review  $\cdot$ computable phenotype  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ residual sampling  $\cdot$ two-phase designs
@@ 1.1 Measuring Whole-Person Health
The advent of learning health systems [1, 2] signals a transition from a reactive (treatment-based) to a proactive (prevention-based) philosophy in healthcare. This transition could prevent future healthcare utilization through early intervention, improving health for the individual and providing cost savings to the institution. Having a standardized summary measure of health that is simple to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
Numerous data types exist within the electronic health record (EHR), including vital signs, demographics, diagnoses, procedures, and clinical notes. Easy access to a standardized and valid summary measure of whole-person health in the EHR would be advantageous for patients, clinicians, and researchers [3, 4]. The electronic frailty index (EFI) serves this purpose but is only designed for those aged 65 and older [5]. A whole-person health metric for adults aged $18+$ would enable system-wide longitudinal tracking of individual and population health, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
Individual physiological variables measure aspects of health; they do not provide a comprehensive assessment across various organ systems. The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [6, 7, 8, 9]. It has been used to predict numerous health outcomes (e.g., [10, 11, 12, 13]) and measure racial disparities in health [14]. Because it is sensitive to short-term interventions, ALI is also an appealing “surrogate” outcome to test health interventions.
Many ALI formulations exist. We focus on the calculation from [7], which computes ALI from ten components across three body systems (Table 1). The original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically-driven thresholds to define binary indicators of being at an unhealthy level (e.g., $=1$ if HbA1c $\geq 6.5$ and $=0$ if HbA1c $<6.5$). ALI is ultimately calculated as the sum of these ten components, with zero and ten representing the best and worst whole-person health, respectively (Supplemental Figure S1).
Whole-person health is a latent concept influenced by genetics, environment, and other factors. Scores like the ALI incorporate components across body systems and yield “computable phenotypes” for it, concisely characterizing patients’ current health status with one-number summaries of observed data [15, 16]. The ALI’s ability to quantify whole-person health has been rigorously examined in controlled studies [6, 7, 8]. However, before adopting the ALI in other settings, it must be revalidated, accounting for real-world data challenges.
@@ 1.2 Data Quality Challenges in Electronic Health Records Data
EHR data present a substantial opportunity to operationalize the ALI in the learning health system. The volume of routinely collected data in the EHR is steadily climbing, increasing the accessibility to clinically meaningful variables. Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains (e.g., [17, 18, 19, 20]).
@@
However, despite clear advantages, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors. Quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [21, 22, 23, 24]. In calculating the ALI, measurements may be (i) missing, making the component missing, or (ii) non-missing but mismeasured, potentially misclassifying the component.
@@
Consider a patient whose true HbA1c was $6.5\%$, so their corresponding component should equal one. We illustrate two scenarios for this patient’s data in Figure 1. In the first, HbA1c was measured but erroneously transposed to $5.6\%$. Based on the EHR, their HbA1c component would be misclassified as zero since $5.6\%$ falls into the healthy range. In the second, HbA1c was not measured, so the value is missing, and we cannot define a component. Counting missing components as zeros would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining ALI as the proportion of non-missing components a patient experienced.
@@
Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation (chart review). In the top scenario, the patient had HbA1c measured but with error, and the correct value was recovered through validation. In the bottom scenario, the patient did not have HbA1c measured, but auxiliary information was located through validation to replace the missing component.
@@
Ignoring errors and missingness in the EHR can hinder analyses, as biased estimates may lead to invalid inference and incorrect conclusions [25, 26]. Errors in EHR data can also reduce the accuracy of prediction models [27]. There is a particular concern for missingness and errors in derived variables, like ALI, that combine data from many different variables [28, 29]. Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about its association with healthcare utilization.
Applying popular missing data approaches like multiple imputation (MI) may offer some improvement at baseline [30]. However, they can also hinder results if any imputation models are misspecified [31]. Since components are likely missing not at random in EHR data (e.g., clinicians usually order labs when they suspect values are abnormal) [32], correctly specifying imputation models could be especially challenging. Plus, MI would not correct for errors in non-missing ALI components.
@@
For the ALI to be useful in patient care and research, it must be reliably and accurately measured. The original calculations, like [7], were done in clinical trials, where all patients must have all measurements, and data validation protocols are likely in place to ensure quality. Similarly, validation via chart review, comparing extracted EHR data to source documents (like electronic charts), provides data quality insights. Complete data validation (for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
@@
Fortunately, partially validating EHR data can help identify issues and pave the way for corrections. It provides a cost-effective alternative to complete data validation and is a promising approach to gauge the quality of observational data sources [33, 34, 35]. These partial validation studies are a special case of two-phase designs [36, 37]. Phase I includes EHR data for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation (chart reviews) in Phase II.
@@
Validated data are usually expected to be more correct (i.e., less error-prone) but not more complete (i.e., less missingness). To calculate the ALI from EHR data, we needed validation to do both. We devise a novel validation protocol that incorporates a “roadmap” created by clinical experts to recover missing data using auxiliary information (e.g., related diagnoses). With it, we verify the ALI’s strength as a computable phenotype for whole-person health and a predictor of healthcare utilization in observational settings with data challenges.
@@
Table 1: Ten components of the allostatic load index were defined by discretizing measurements across three body systems at clinically-driven thresholds. If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic (the institution’s electronic charting software). If the auxiliary information was present, the patient was treated as having an “unhealthy” measurement and the component included in their validated ALI.
1.3 Application to the Learning Health System
We investigate how to overcome data challenges to measure whole-person health across one of the largest learning health systems in the United States. Having recently integrated the EFI into the EHR to measure whole-person health for patients $65$ and older [38], the health system is interested in operationalizing the ALI for the broader patient population. We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina. Our sample was drawn from patients (i) $18$-$65$ years old who (ii) initiated their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic). Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
Our data contained demographics, vitals, lab values, discharge diagnoses, and problem lists dating $3$ years before the first encounter. Key variables included measurements for the ten ALI components (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period. Since most patients had multiple encounters, ALI components were based on mean measurements.
@@ 1.4 Overview
While very insightful in the development and use of computable phenotypes, partial validation studies are expensive, time-intensive undertakings. Using a representative sample of $1000$ patients from the EHR at an extensive academic learning health system ($100$ of whom could be validated), we propose novel methods to design, conduct, and analyze statistically efficient and robust studies of whole-person health (ALI) and healthcare utilization. Using advanced statistics and bioinformatics methods, we seize three key opportunities to maximize our investment into partial validation: robust and efficient statistical modeling, enriched chart review protocol, and targeted sampling design.
@@
  • Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in electronic health records (EHR) to make patient- and institution-level decisions. However, routinely collected data are prone to missingness and errors.
@@
  • Who Would Benefit From the Knowledge in This Paper Biomedical researchers interested in analyzing routinely collected data, like those in the EHR, while correcting their data quality issues.
@@ 2 Methods
2.1 Healthcare Utilization Model
This study was approved by the Institutional Review Board at Wake Forest University School of Medicine (WFUSM). We model healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period. Whole-person health, measured by ALI ($X\in[0,1]$), is the exposure of interest. Each component $S_{j}$ ($j\in\{1,\dots,10\}$) equals one if the patient’s measurement was considered unhealthy and zero otherwise, such that $X=\sum_{j=1}^{10}S_{j}/10$ is the proportion of unhealthy measurements experienced. Further, we control for age ($Z\in\mathcal{R}$).
We assume a logistic regression model for the log-odds that $Y=1$ (patient engaged in healthcare system) given ($X$, $Z$). A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (logORs), denoted ${\boldsymbol{\beta}}$. Age $Z$ is fully observed and expected to be error-free. However, due to missingness and errors in ALI components, only an error-prone measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available from the EHR data. We define error-prone components $S_{j}^{*}$ ($j\in\{1,\dots,10\}$) and introduce missingness indicators $M_{j}^{*}$ for them (i.e., $M_{j}^{*}=1$ if $S_{j}^{*}$ is missing). We compute $X^{*}=\sum_{j=1}^{10}(1-M_{j}^{*})S_{j}^{*}/\sum_{j=1}^{10}(1-M_{j}^{*})$, i.e., as the proportion of non-missing components experienced. Calculating ALI as a proportion – rather than sum – was critical. Summing over non-missing components treats missing values as zeros (i.e., healthy measurements); instead, the proportion ignores them.
2.2 Three Key Steps to Partial Validation Studies: Analysis, Design, and Protocol
To obtain consistent estimates for the healthcare utilization model, we planned a partial validation study. We first extracted EHR data ($Y$, $X^{*}$, $Z$) for a sample of $N=1000$ patients. Then, a subset of $n=100$ patients was selected for validation via chart review such that $X$ was also measured.
Partial validation studies break down into three key steps. Design: Select which patients should be validated. Protocol: Develop standardized tools and practices for those patients’ chart reviews. Analysis: Estimate statistical models or quantities of interest from validated and unvalidated data together. These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information. For added flexibility, a multi-wave validation framework, conducting chart reviews in multiple sprints, allows us to revisit key steps and make modifications as information accumulates (Figure 2).
In our application (Section 3), we built the statistical model and developed the baseline sampling strategy and tools before beginning validation. The proposed methods are detailed in order of development: analysis (Section 2.3), design (Section 2.4), and protocol (Section 2.5). Across three validation waves, we updated our sampling strategy and tools (revisiting the design and protocol steps, respectively), before fitting the final model (Section 2.6).
Figure 2: After extracting the electronic health records (EHR) data, partial validation studies break down into three key steps: analysis, design, and protocol. These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information. For added flexibility, a multi-wave validation framework allows us to revisit these key steps and make modifications as information accumulates.
2.3 Analysis: Robust and Efficient Statistical Model for Partially Validated Data
Let $V_{i}$ distinguish validated and unvalidated patients, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise. Incorporating EHR and chart review data, along with the validation indicator, the joint distribution of a complete observation is assumed to be
@@ 2.3 Analysis: Robust and Efficient Statistical Model for Partially Validated Data
where $\Pr(V\mid Y,X^{*},Z)$ is the validation sampling probability; ${\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)$ is the logistic regression model; $\Pr(X\mid X^{*},Z)$ is the exposure error mechanism; and $\Pr(X^{*},Z)$ is the joint probability density or mass function of ($X^{*}$, $Z$). Two assumptions are made in (1). First, selection for validation depends only on fully observed variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [39]. Second, $X^{*}$ from the EHR is a surrogate for validated $X$ from the chart reviews, such that $Y$ and $X^{*}$ are conditionally independent given $X$.
Observations for all $N=1000$ patients are assumed to be independent and identically distributed following (1). Thus, the observed-data log-likelihood function (hereafter, log-likelihood) for the logORs ${\boldsymbol{\beta}}$ is defined from (1), after two simplifications. First, we can omit $\Pr(V|Y,X^{*},Z)$ from the log-likelihood because $X$ is MAR. Second, we can ignore $\Pr(X^{*},Z)$ because it drops out as the log-likelihood is maximized with respect to ${\boldsymbol{\beta}}$ since ($X^{*}$, $Z$) are fully observed. Now, the log-likelihood for ${\boldsymbol{\beta}}$ is proportional to
@@
Notice that unvalidated patients contribute incomplete information to (2) by summing over the missing validated $X$:
@@
A model for $\Pr(X\mid X^{*},Z)$ is needed to estimate ${\boldsymbol{\beta}}$ from (2) since $X$ is only observed for validated patients. Since $X\in[0,1]$, there are not many flexible, parametric choices to model its distribution given ($X^{*},Z$). A nonparametric estimator is preferable, allowing $X$ to take on any values and arbitrarily depend on ($X^{*}$, $Z$).
@@
[40] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described. The current setting is a special case with only exposure error, and a brief outline is provided in the Supplemental Materials. The logiSieve package contains R functions to fit this model [41]. For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see [40].
2.4 Design: Targeted Sampling to Select Informative Patients for Validation